OpenStudy (anonymous):

fan and mental what is 1/2x(13/6+7/8)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Is this it? \(\dfrac{1}{2}x(\dfrac{13}{6} + \dfrac{7}{8})\)

OpenStudy (anonymous):

a 29/48 b 1 1/24 c 1 25/48 d 2 25/48

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

times

OpenStudy (anonymous):

(13/​6+​7/​8)

OpenStudy (igreen):

First simplify the parenthesis: \(\dfrac{1}{2}(3 \dfrac{1}{24})\) Now distribute 1/2: \(1 \dfrac{25}{48}\)

OpenStudy (anonymous):

ty

OpenStudy (igreen):

Understand? @Pla

OpenStudy (anonymous):

wait

OpenStudy (igreen):

@PlanSlam

OpenStudy (igreen):

Yes?

OpenStudy (anonymous):

i get it ty

OpenStudy (igreen):

Okay, no problem.